Grade 11General Physics

If a ,b and c are unit vectors such that a +b -c = 0 ,then the angle between a and b

Prachi
10 Years agoGrade 11

1 Answer

Vikas TU
ApprovedApproved Tutor Answer10 Years ago
cos@  = a.b/|a||b| = a.b {magnitude of abc vectors being 1 as thy r unit vectors}
cos@ = a.b
Noe from
a+b-c = 0
a+b =c
sq. both sides,
|a|^2 + |b|^2 + 2a.b = |c|^2
1 + 1 + 2a.b = 1
a.b = -1/2
 
therefore,
cos@ = -1/2
@ = 120 degree..